# ShowTable: Unlocking Creative Table Visualization with Collaborative Reflection and Refinement

### v1

@@ 1 Introduction
Figure 1: The illustration of our proposed creative table visualization task and ShowTable pipeline. Given a table about a specific topic, our task requires the model to produce a visualization infographic that is aesthetic and faithful to the data points. The ShowTable pipeline employs four steps: rewriting, generation, reflection, and refinement, thus achieving high-fidelity visualization. We use Wan2.5-Preview [wan2.5] here for generation and refinement.
Image generation models have shown significant growth in quality in recent years [sd, sd3, dalle3, flux, t2i_r1]. Recent advancements, particularly in unified models [januspro, showo, emu3, blip3o, bagel, uniworld, omnigen2, qwenimage, wan2.5], have leveraged MLLMs for fine-grained understanding of complex multi-modal instructions, leading to more semantically coherent synthesis in general-purpose image generation and editing. Despite this rapid progress in the general domain, research is increasingly shifting towards more complex scenarios. A key area of focus is graphic design, such as poster generation [autoposter, postermaker, dreamposter] and text rendering [tuo2024anytext2, liu2024glyph, chen2024textdiffuser, qwenimage]. These tasks present a higher challenge as they demand not photorealism, but rather precise textual alignment and adherence to aesthetic principles of graphic design.
@@ 2 Related Work
Image generation for graphic design. Recent advances in image generation have significantly enhanced the quality and semantic controllability of synthesized content in general-purpose scenarios [sd, sd3, dalle3, flux, t2i_r1, januspro, showo, emu3, blip3o, bagel, uniworld, omnigen2, qwenimage, wan2.5]. This progress has spurred growing interest in more complex generation tasks such as graphic design. AnyText [tuo2023anytext, tuo2024anytext2], Glyph-ByT5 [liu2024glyph], and TextDiffuser [chen2024textdiffuser] have made significant efforts in text rendering, focusing on accurate textual element incorporation. AutoPoster [autoposter], PosterMaker [postermaker], and DreamPoster [dreamposter] focus on poster generation, which addresses aesthetic layout planning. The recently proposed Qwen-Image [qwenimage] has also demonstrated capabilities in complex text rendering and infographic design. Compared to these scenarios, our proposed creative table visualization task presents a more formidable challenge. Our task demands not only reasoning about graphic design for aesthetic layout but also a precise data-to-visual mapping for faithful representation of tabular content.
Reasoning and reflection paradigm of MLLMs. The reasoning and reflection abilities of MLLMs are gaining increasing attention for both understanding and generation. Some studies leverage the reasoning capability of MLLMs to achieve better image understanding, often termed thinking with images [thinkingwithgeneratedimages, zheng2025deepeyes, zhuo2024lumina, Visualsketchpad]. For image generation, some works [bagel, t2i_r1, fang2025got, xiao2025mindomni] introduce a text-based reasoning step, enhancing image generation performance. Recently, emerging research has begun to leverage the MLLM reflection process to refine the image generation itself, aiming to enhance instruction-following in complex, general-purpose scenes [IRG, huang2024dialoggen, wang2024genartist]. However, this category of research remains focused on general domains. The high information density and specific structural constraints of our task pose a greater challenge for MLLMs with combining reasoning, planning, and reflection abilities.
@@ 3.2 Training Details
Rewriting module. Although general LLMs have demonstrated powerful reasoning abilities on various tasks, their performance on table reasoning and compositional planning can be further enhanced. As Figure 3 shows, they may miss data points or plan poorly when encountering complex data (e.g., multi-layered structures). As this is the critical step responsible for basic layout design and content planning, fine-tuning the rewrite module is expected to significantly enhance the overall performance. Therefore, we fine-tune a specific rewrite module from Qwen3-8B [qwen3] to address this challenge. With our constructed 30K rewriting training data (Section 4.2), we train the module with the standard next-token prediction format:
@@
Refinement module. The refinement stage is critical for correcting errors, but it may also present a significant potential bottleneck. As illustrated in Figure 4, our preliminary experiments revealed this challenge. When employing a recent SOTA Qwen-Image-Edit [qwenimage] within our iterative loop, we observed a performance degradation with each correction round. This raised a critical question: is the pipeline’s self-correction logic flawed, or is the editing model’s capability insufficient? To investigate this variable, we introduced Wan2.5-I2I-Preview [wan2.5], another editing model known for fine-grained controllability, and found that performance indeed increased with each iteration. This indicates that our pipeline structure is sound, but its effectiveness is fundamentally constrained by the refinement model’s ability to execute precise edits. Therefore, to resolve this bottleneck, we train the refinement module using RL, specifically employing the Group Relative Policy Optimization (GRPO) algorithm [deepseekr1]. This approach requires an accurate reward signal. However, evaluating rendering quality is complex, integrating multiple dimensions. As recent pretrained MLLMs struggle to directly provide consistent, accurate scalar scores for such assessments [luo2025editscore] (see Appendix), there remains a need to develop a specialized reward model (RM). Using our constructed 30K pairwise preference dataset (Section 4.2), we fine-tune a Qwen2.5-VL-3B [qwen2.5vl] model, $f_{\theta}$, as our quality assessor. The model is trained to distinguish positive ($x_{w}$) and negative ($x_{l}$) samples for a given prompt $p$ using the Bradley-Terry loss [bradley1952bradterry]:
$\mathcal{L}_{BT}=-\mathbb{E}_{(p,x_{w},x_{l})\sim D}\left[\log\sigma\left(f_{\theta}(x_{w},p)-f_{\theta}(x_{l},p)\right)\right],$ (2)
where $\sigma(\cdot)$ is the Sigmoid function. To enhance training efficiency, the scalar reward score is computed by extracting the probabilities corresponding to digits 0–9 from the output logits and averaging their sum. Finally, using our constructed 5K refinement data (Section 4.2), we follow Flow-GRPO [flowgrpo] to perform the RL. The full reward signal combines our trained $f_{\theta}$ with an existing aesthetic reward model, ImageReward [imagereward], to optimize the training.
@@ 4.1 Data Collection and Filtering
To support both training and benchmarking for our proposed task, we construct a dataset of 30K high-quality table-image pairs and an additional benchmark of 800 evaluation samples. We initiate this by collecting raw images from diverse public datasets, including SlideVQA [tanaka2023slidevqa], OpenImages [kuznetsova2020open], and Cambrian-10M [tong2024cambrian]. We then applied the rigorous filtering and annotation pipeline shown in Figure 5. First, we discard low-resolution images (under $200\times 200$) and images lacking text, as identified by PaddleOCR [cui2025paddleocr]. Following this, we utilize SOTA MLLMs (i.e., Gemini-2.5-pro[gemini2.5pro] and GPT-5 [gpt5]) to perform a crucial filtering and annotation step: the models filter out images that do not feature statistical data as the main body, and simultaneously annotate the table information in markdown format for all remaining images. To ensure the quality of these annotations, we implemented a consensus-based verification process. Both MLLMs independently annotate the filtered images. We then retained only the samples for which the annotations from both models were consistent and mutually approved, resulting in our final set of 30K high-quality table-image pairs.
@@ 4.2 Training Data Construction
Refinement training data. To generate data for RL, we first use the descriptions from the previous step to generate initial images using Wan2.5-t2i-preview [wan2.5] and Qwen-Image [qwenimage]. We then use Gemini-2.5-pro to audit these generated images and produce corresponding refinement instructions. We apply a rigorous filtering strategy to ensure RL training stability. For each sample (initial image + instruction), we instruct our base editing model (Qwen-Image-Edit [qwenimage]) to generate five refined candidates. These candidates are compared against the initial image by a powerful MLLM assessor (GPT-5). We discard samples where all five attempts are judged as worse, or all five as better, than the original. This filtering isolates samples that are too hard or too easy for base model, yielding 5K challenging samples ideal for refinement training.
Reward training data. Our RL approach requires a reliable RM. Given that MLLMs are unstable in providing direct point-wise scores [luo2025editscore], we construct a preference dataset for converting the preference into a specific score assessor. We use GPT-5 and Gemini-2.5-pro to compare each image pair, followed by voting, and finally generate 30K positive-negative image pairs for training the RM, as shown in Figure 5. These pairs are sourced from three comparisons: 1) images refined by Wan2.5 versus images generated by Wan2.5 or Qwen; 2) images generated by Wan2.5 versus those by Qwen; and 3) collected ground-truth images versus generated or refined images.
@@
Blip3o-Next [chen2025blip3onext] 0.4 18.0 4.4 6.2 2.5 10.8
@@ 4.3 Benchmark Design
For comprehensive benchmarking, we conduct multi-view evaluation [liu2025capability] with five well-designed dimensions, focusing not only on factual accuracy but also on logical coherence and visual aesthetics. Instead of using an MLLM as a subjective scorer, we leverage it as a quality assurance analyst. For the first four dimensions, the MLLM is prompted to identify and quantify specific errors within the generated chart. The final score is then deterministically calculated based on the number of reported errors, thereby mitigating the instability and bias associated with direct LLM-based scoring. The final dimension is quantitatively assessed using a dedicated aesthetic scoring model. The five dimensions are as follows:
@@
Qwen3-8B [qwen3] 30.6 71.5 46.6 34.1 5.1 46.8
GPT-5 [gpt5] 35.9 78.5 47.8 41.8 5.2 51.2
Gemini-2.5-pro [gemini2.5pro] 40.8 79.9 53.9 41.1 5.1 53.3
@@ 5.1 Implementation Setup
The modular design of our ShowTable pipeline allows for flexible combinations of MLLMs and diffusion models. In our default configuration, the rewriting module is trained based on Qwen3-8B [qwen3], and the refinement module is trained based on a distilled 8-step version [qwenimagelightning] of Qwen-Image-Edit-2509 [qwenimage] to accelerate RL training. For the reflection module, we employ GPT-5-2025-08-07 [gpt5]. For fair comparison, all baselines that use Qwen-Image for generation or Qwen-Image-Edit-2509 for refinement also use these same distilled versions [qwenimagelightning]. We set the maximum self-correction round to 3, though the process can terminate early if the reflection module deems an image satisfactory. More detailed settings can be found in the Appendix.
@@ 5.2 Main Results
We comprehensively evaluate the effectiveness of the ShowTable pipeline by applying it to several advanced T2I generation models, including Flux [flux], Bagel [bagel], Blip3o-Next [chen2025blip3onext], UniWorld-V1 [uniworld], OmniGen2 [omnigen2], and Qwen-Image [qwenimage]. The evaluation is conducted on our TableVisBench based on the five dimensions. We systematically compare performance under three configurations: 1) The base generation model alone (Base); 2) The base model prefixed with our rewriting module (RW+Base); 3) The full pipeline, integrating all modules (RW+Base+REF).
@@
Qwen3-VL-235B [qwen3vl] 37.7 78.1 45.1 31.3 4.4 47.2
Gemini-2.5-pro [gemini2.5pro] 43.2 81.2 46.6 41.1 4.4 51.2
GPT-5 [gpt5] 42.6 79.7 45.0 35.9 4.4 49.4
@@
Qwen3-VL-235B [qwen3vl] 46.9 81.5 48.3 37.1 4.5 51.8
Gemini-2.5-pro [gemini2.5pro] 48.1 83.0 48.8 44.8 4.5 53.9
GPT-5  [gpt5] 52.4 82.9 54.3 40.0 4.5 54.9
5.3 Ablation Studies
@@ 6 Conclusion
Appendix
@@ A.1 Chart Generation with Agentic Tools
With rapid development of current MLLMs [llava, qwen2.5vl, liu2025hybrid, zheng2025deepeyes], the strong tool calling ability provides a promising way to complete multi-modal tasks. Current approaches for chart generation predominantly rely on LLMs coupled with agentic tools, generally falling into three main categories. The first category [tian2024chartgpt, rashid2022text2chart, zhang2024chartifytext] employs LLMs in an end-to-end manner to produce charts directly from text. These methods typically parse input to identify axes, map entities, and classify chart types, subsequently generating structured specifications for rendering engines like Vega-Lite. While automated, their expressiveness is strictly confined by the predefined grammatical rules of the underlying visualization language. The second category [zhang2025thyme, zhao2025pyvision, zhao2025chartedit, zadeh2024text2chart31, han2023chartllama] focuses on generating executable plotting code (e.g., via Matplotlib/Python). Recent works have incorporated multi-agent frameworks and reflection mechanisms to improve the syntactic correctness of the generated code [ford2025does]. Although these methods offer precise control, they heavily depend on external rendering engines and typically lack the capability to handle complex, artistic visual designs beyond standard plots. The third category [xiao2023let] adopts a retrieval-editing pipeline that selects visual templates from an image corpus and adapts them to new data. While this benefits from template reuse, it is limited by the diversity of the database and often faces challenges in accurately aligning new data with retrieved visual structures.
Discussion. While existing methods excel in structural automation and factual plotting, they fundamentally struggle with creativity and aesthetics. Code-based and template-based approaches are bound by rigid rendering logic, making it difficult to produce visually striking infographics suitable for professional poster design, slide generation, or data-driven storytelling. They prioritize “correctnes” but often fail at “presentation.” In contrast, our work proposes the Creative Table Visualization task to explore the untapped potential of generative models and unified models in this domain. We argue that generative models offer a significantly higher ceiling for flexibility and aesthetic quality, capable of seamlessly integrating data into artistic compositions. By validating the feasibility of this approach, we aim to break the traditional boundaries of rule-based rendering and pave the way for more robust, unified, and creative visual synthesis systems.

### v2

@@ 1 Introduction
Figure 1: The illustration of our proposed creative table visualization task and ShowTable pipeline. Given a table about a specific topic, our task requires the model to produce a visualization infographic that is aesthetic and faithful to the data points. The ShowTable pipeline employs four steps: rewriting, generation, reflection, and refinement, thus achieving high-fidelity visualization. We use Wan2.5-Preview [51] here for generation and refinement.
Image generation models have shown significant growth in quality in recent years [46, 15, 43, 5, 28, 63]. Recent advancements, particularly in unified models [10, 70, 58, 8, 14, 34, 66, 50, 65, 51], have leveraged MLLMs for fine-grained understanding of complex multi-modal instructions [39], leading to more semantically coherent synthesis in general-purpose image generation and editing. Despite this rapid progress in the general domain, research is increasingly shifting towards more complex scenarios such as graphic design (e.g., poster generation [35, 19, 24] and text rendering [55, 38, 7]). These tasks present a higher challenge as they demand not photorealism, but precise textual alignment and adherence to aesthetic principles of graphics.
@@ 2 Related Work
Image generation for graphic design. Recent advances in image generation have significantly enhanced the quality and semantic controllability of synthesized content in general-purpose scenarios [46, 15, 43, 5, 28, 10, 70, 58, 8, 14, 34, 66, 65, 51]. This progress has spurred growing interest in more complex generation tasks such as graphic design. AnyText [56, 55], Glyph-ByT5 [38], and TextDiffuser [7] have made significant efforts in text rendering, focusing on accurate textual element incorporation. AutoPoster [35], PosterMaker [19], and DreamPoster [24] focus on poster generation, which addresses aesthetic layout planning. The recently proposed Qwen-Image [65] also emphasizes capabilities in complex text rendering and infographic design. Compared to these scenarios, our proposed creative table visualization task presents a more formidable challenge. Our task demands not only reasoning about graphic design for aesthetic layout but also a precise data-to-visual mapping for faithful representation of tabular content.
Reasoning and reflection paradigm of MLLMs. The reasoning and reflection abilities of MLLMs are gaining increasing attention for both understanding and generation. Some studies leverage the reasoning capability of MLLMs to achieve better image understanding, often termed thinking with images [11, 80, 81, 25, 3]. For image generation, some works [14, 28, 17, 69] introduce a text-based reasoning step, enhancing image generation performance. Recently, emerging research has begun to leverage the MLLM reflection process to refine the image generation itself, aiming to enhance instruction-following in complex, general-purpose scenes [27, 26, 60]. However, this category of research remains focused on general domains. The high information density and specific structural constraints of our task pose a greater challenge for MLLMs with combining reasoning, planning, and reflection abilities.
@@ 3.2 Training Details
Rewriting module. Although general LLMs have demonstrated powerful reasoning abilities on various tasks, their performance on table reasoning and compositional planning can be further enhanced. As Figure 3 shows, they may miss data points or plan poorly when encountering complex data (e.g., multi-layered structures). As this is the critical step responsible for basic layout design and content planning, fine-tuning the rewrite module is expected to significantly enhance the overall performance. Therefore, we fine-tune a specific rewrite module from Qwen3-8B [73] to address this challenge. With our constructed 30K rewriting training data (Section 4.2), we train the module with the standard next-token prediction format:
@@
Refinement module. The refinement stage is critical for correcting errors, but it may also present a significant potential bottleneck. As illustrated in Figure 4, our preliminary experiments revealed this challenge. When employing a recent SOTA Qwen-Image-Edit [65] within our iterative loop, we observed a performance degradation with each correction round. This raised a critical question: is the pipeline’s self-correction logic flawed, or is the editing model’s capability insufficient? To investigate this variable, we introduced Wan2.5-I2I-Preview [51], another editing model known for fine-grained controllability, and found that performance indeed increased with each iteration. This indicates that our pipeline structure is sound, but its effectiveness is fundamentally constrained by the refinement model’s ability to execute precise edits. Therefore, to resolve this bottleneck, we train the refinement module using RL, specifically employing the Group Relative Policy Optimization (GRPO) algorithm [21]. This approach requires an accurate reward signal. However, evaluating rendering quality is complex, integrating multiple dimensions. As recent pretrained MLLMs struggle to directly provide consistent, accurate scalar scores for such assessments [41] (see Appendix), there remains a need to develop a specialized reward model (RM). Using our constructed 30K pairwise preference dataset (Section 4.2), we fine-tune a Qwen2.5-VL-3B [2] model, $f_{\theta}$, as our quality assessor. The model is trained to distinguish positive ($x_{w}$) and negative ($x_{l}$) samples for a given prompt $p$ using the Bradley-Terry loss [6]:
$\mathcal{L}_{BT}=-\mathbb{E}\left[\log\sigma\left(f_{\theta}(x_{w},p)-f_{\theta}(x_{l},p)\right)\right],$ (2)
where $\sigma(\cdot)$ is the Sigmoid function. To enhance training efficiency, the scalar reward score is computed by extracting the probabilities corresponding to digits 0–9 from the output logits and averaging their sum. Finally, using our constructed 5K refinement data (Section 4.2), we follow Flow-GRPO [37] to perform the RL. The full reward signal combines our trained $f_{\theta}$ with an existing aesthetic reward model, ImageReward [71], to optimize the training.
@@ 4.1 Data Collection and Filtering
To support both training and benchmarking for our proposed task, we construct a dataset of 30K high-quality table-image pairs and an additional benchmark of 800 evaluation samples. We initiate this by collecting raw images from diverse public datasets, including SlideVQA [49], OpenImages [31], and Cambrian-10M [54]. We then applied the rigorous filtering and annotation pipeline shown in Figure 5. First, we discard low-resolution images (under $200\times 200$) and images lacking text, as identified by PaddleOCR [13]. Following this, we utilize SOTA MLLMs (i.e., Gemini-2.5-pro[12] and GPT-5 [48]) to perform a crucial filtering and annotation step: the models filter out images that do not feature statistical data as the main body, and simultaneously annotate the table information in markdown format for all remaining images. To ensure the quality of these annotations, we implemented a consensus-based verification process. Both MLLMs independently annotate the filtered images. We then retained only the samples for which the annotations from both models were consistent and mutually approved, resulting in our final set of 30K high-quality table-image pairs.
@@ 4.2 Training Data Construction
Refinement training data. To generate data for RL, we first use the descriptions from the previous step to generate initial images using Wan2.5-t2i-preview [51] and Qwen-Image [65]. We then use Gemini-2.5-pro to audit these generated images and produce corresponding refinement instructions. We apply a rigorous filtering strategy to ensure RL training stability. For each sample (initial image + instruction), we instruct our base editing model (Qwen-Image-Edit [65]) to generate five refined candidates. These candidates are compared against the initial image by a powerful MLLM assessor (GPT-5). We discard samples where all five attempts are judged as worse, or all five as better, than the original. This filtering isolates samples that are too hard or too easy for base model, yielding 5K challenging samples ideal for refinement training.
Reward training data. Our RL approach requires a reliable RM. Given that MLLMs are unstable in providing direct point-wise scores [41], we construct a preference dataset for converting the preference into a specific score assessor. We use GPT-5 and Gemini-2.5-pro to compare each image pair, followed by voting, and finally generate 30K positive-negative image pairs for training the RM, as shown in Figure 5. These pairs are sourced from three comparisons: 1) images refined by Wan2.5 versus images generated by Wan2.5 or Qwen; 2) images generated by Wan2.5 versus those by Qwen; and 3) collected ground-truth images versus generated or refined images.
@@
Blip3o-Next [9] 0.4 18.0 4.4 6.2 2.5 10.8
@@ 4.3 Benchmark Design
For comprehensive benchmarking, we conduct multi-view evaluation [40] with five well-designed dimensions, focusing not only on factual accuracy but also on logical coherence and visual aesthetics. Instead of using an MLLM as a subjective scorer, we leverage it as a quality assurance analyst. For the first four dimensions, the MLLM is prompted to identify and quantify specific errors within the generated chart. The final score is then deterministically calculated based on the number of reported errors, thereby mitigating the instability and bias associated with direct LLM-based scoring. The final dimension is quantitatively assessed using a dedicated aesthetic scoring model. The five dimensions are as follows:
@@
Qwen3-8B [73] 30.6 71.5 46.6 34.1 5.1 46.8
GPT-5 [48] 35.9 78.5 47.8 41.8 5.2 51.2
Gemini-2.5-pro [12] 40.8 79.9 53.9 41.1 5.1 53.3
@@ 5.1 Implementation Setup
The modular design of our ShowTable pipeline allows for flexible combinations of MLLMs and diffusion models. In our default configuration, the rewriting module is trained based on Qwen3-8B [73], and the refinement module is trained based on a distilled 8-step version [42] of Qwen-Image-Edit-2509 [65] to accelerate RL training. For the reflection module, we employ GPT-5-2025-08-07 [48]. For fair comparison, all baselines that use Qwen-Image for generation or Qwen-Image-Edit-2509 for refinement also use these same distilled versions [42]. We set the maximum self-correction round to 3, though the process can terminate early if the reflection module deems an image satisfactory. More detailed settings can be found in the Appendix.
@@ 5.2 Main Results
We comprehensively evaluate the effectiveness of the ShowTable pipeline by applying it to several advanced T2I generation models, including Flux [5], Bagel [14], Blip3o-Next [9], UniWorld-V1 [34], OmniGen2 [66], and Qwen-Image [65]. The evaluation is conducted on our TableVisBench based on the five dimensions. We systematically compare performance under three configurations: 1) The base generation model alone (Base); 2) The base model prefixed with our rewriting module (RW+Base); 3) The full pipeline, integrating all modules (RW+Base+REF).
@@
Qwen3-VL-235B [1] 37.7 78.1 45.1 31.3 4.4 47.2
Gemini-2.5-pro [12] 43.2 81.2 46.6 41.1 4.4 51.2
GPT-5 [48] 42.6 79.7 45.0 35.9 4.4 49.4
@@
Qwen3-VL-235B [1] 46.9 81.5 48.3 37.1 4.5 51.8
Gemini-2.5-pro [12] 48.1 83.0 48.8 44.8 4.5 53.9
GPT-5  [48] 52.4 82.9 54.3 40.0 4.5 54.9
@@ 6 Conclusion
Acknowledgment
This work is supported by the National Nature Science Foundation of China (62425114, 62121002, U23B2028), and the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (JYB2025XDXM103). We acknowledge the support of Alibaba Group, the GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC, and USTC super-computing center for providing computational resources for this project.
Supplementary Material
@@ A.1 Chart Generation with Agentic Tools
With rapid development of current MLLMs [36, 2, 39, 80], the strong tool calling ability provides a promising way to complete multi-modal tasks. Current approaches for chart generation predominantly rely on LLMs coupled with agentic tools, generally falling into three main categories. The first category [53, 45, 75] employs LLMs in an end-to-end manner to produce charts directly from text. These methods typically parse input to identify axes, map entities, and classify chart types, subsequently generating structured specifications for rendering engines like Vega-Lite. While automated, their expressiveness is strictly confined by the predefined grammatical rules of the underlying visualization language. The second category [76, 77, 78, 74, 22] focuses on generating executable plotting code (e.g., via Matplotlib/Python). Recent works have incorporated multi-agent frameworks and reflection mechanisms to improve the syntactic correctness of the generated code [18]. Although these methods offer precise control, they heavily depend on external rendering engines and typically lack the capability to handle complex, artistic visual designs beyond standard plots. The third category [68] adopts a retrieval-editing pipeline that selects visual templates from an image corpus and adapts them to new data. While this benefits from template reuse, it is limited by the diversity of the database and often faces challenges in accurately aligning new data with retrieved visual structures.
Table A1: The comparison between our ShowTable pipeline and code-based methods. For our ShowTable pipeline, we use Gemini-2.5-pro for rewriting, Wan2.5-T2I-Preview for generation, GPT-5 for Reflection, and Wan2.5-I2I-Preview for refinement here. For code-based methods, we use Gemini-2.5-pro to write python code with the matplotlib library.
Methods DA TR $R$ AA AQ Score
ShowTable Pipeline 69.6 84.9 67.6 68.4 4.9 67.9
Code (Gemini-2.5-pro) 83.8 89.2 85.3 97.7 4.1 79.4
Figure A1: The qualitative comparison between code-based methods and our ShowTable pipeline. The code-based methods often prioritize “correctness” but may fail at “presentation”
Discussion. While existing methods excel in structural automation and factual plotting, they fundamentally struggle with creativity and aesthetics. Code-based and template-based approaches are bound by rigid rendering logic, making it difficult to produce visually striking infographics suitable for professional poster design, slide generation, or data-driven storytelling. They prioritize “correctness” but often fail at “presentation”. To explicitly validate this, we conducted a quantitative comparison between a strong code-based baseline (e.g., using Gemini-2.5-Pro to generate executable plotting code) and our ShowTable pipeline, as shown in Tab. A1 and  Fig. A1. As expected, code generation methods inherently achieve superior factual accuracy across structural metrics such as Data Accuracy and Relative Relationship. However, they fall significantly short in Aesthetic Quality (AQ: 4.1 vs. our 4.9) and frequently produce visual artifacts, such as overlapping text. This rigidity restricts their utility in design-centric, creative scenarios.
Importantly, our objective is not to replace traditional code-based rendering tools where absolute numerical precision is the sole priority. Instead, our work proposes the Creative Table Visualization task to explore the untapped potential of generative models and unified models in this domain. We frame our contribution as a technical advancement in multi-modal generation control: we aim to retain the unparalleled creative flexibility of generation models while actively mitigating their primary weakness, data fidelity, through our progressive self-correcting pipeline. We argue that generative models offer a significantly higher ceiling for flexibility and aesthetic quality, capable of seamlessly integrating data into artistic compositions. Furthermore, equipping text-to-image models with rigorous data fidelity is a vital frontier for achieving more general capabilities in visual synthesis (including scientific reporting), as seen in the trajectory of recent models like Wan2.6 [52] and Nano Banana [20] after the CVPR submission. By validating the feasibility of this approach, we aim to break the traditional boundaries of rule-based rendering and pave the way for more robust, unified, and creative visual synthesis systems.